OpenStudy (anonymous):

how do u solve log (base5) of 5/y

OpenStudy (anonymous):

is the question \[\log_{5} \frac{ 5 }{ y }\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Is there something on the other side of the equation? Because you can't solve for y without anything on the other side.

OpenStudy (anonymous):

no i am guessing you set it = to 0

OpenStudy (anonymous):

alright then, so one property of logs is that \[\log_{a}\frac{ x }{ z } = \log_{a}x - \log_{a}z\]

OpenStudy (anonymous):

so\[\log _{5}5-\log _{5}y\]

OpenStudy (anonymous):

yup, now just set that equal to zero and solve

OpenStudy (anonymous):

ok but how do u get rid of the base

OpenStudy (anonymous):

that's another property of logs \[\log_{x}x = 1\] and to get rid of the log base you just do \[x^{\log_{x} }\]